OpenStudy (anonymous):

-5y+2(2y+1)=2(5y-1)-7

OpenStudy (anonymous):

y = 1

OpenStudy (anonymous):

-5y +4y + 2 = 10y -2 -7 -1y + 2 = 10y - 9 -11y = -11 y = 1

OpenStudy (anonymous):

correct answer